OpenStudy (anonymous):

How do I solve for a? 3x-5a/a=y

OpenStudy (anonymous):

\[\frac{5x-a}{a}=y\] like that?

OpenStudy (anonymous):

3x-5a over a

OpenStudy (anonymous):

\[\frac{3x-5a}{a}=y\]

OpenStudy (anonymous):

that one?

OpenStudy (anonymous):

\[\frac{ 3x - 5a }{ a }=y\] \[\frac{ 3x }{ a } - \frac{5a}{a} = y\] The 5a/a becomes 5 because a/a is cancelled out. \[\frac{ 3x-5 }{ a }=y\] \[3x-5=ya\] \[\frac{ 3x-5 }{ y }=a\]

OpenStudy (anonymous):

Sorry, I made a mistake in the equation. The correct solution after a/a is cancelled out from 5a/a is: \[\frac{ 3x }{a}-5=y\] \[\frac{ 3x }{ a }=y+5\] which is the same as \[\frac{ 3x }{ y+5 } = a\]